This question will involve consideration of conditions observed on two different days. On Day One, the surface temperature is 25°C and the surface dewpoint is 15°C. On Day Two, the surface temperature is 25°C and the surface dewpoint temperature is 20°C. If cumulus clouds form on both days, on which of the two days is the cloudbase (i.e. the height of the bottom of the cloud) lower? Clearly explain the reasoning you used to arrive at your answer
Answer #1

Cloud base defines the lowest altitude at which the clouds are formed. Clouds are formed when the air gets saturated with the water vapour present in it. At this situation air temperature become equal to the dew point temperature. For dry adiabatic rate air temperature decreases at a rate of 10℃/1000m and the dew point temperature decreases at a rate of 2℃/1000m. That means the difference between the dew point temperature and the air temperature approaches each other at a rate of 8℃/1000m.

So, the height of the cloud base can be determined if we know the surface temperature and dew point.

For the first case height of the cloud base=125(25 - 15) m=1250 m

For the second case height of the cloud base=125(25 - 20) m=625 m

So, the height of the bottom of the cloud is lower for the second place.